# Non-stationary Gaussian random fields on hypersurfaces: Sampling and strong error analysis

### v1

@@ B.1. Geometric consistency estimate
The following geometric consistency estimate quantifies the error between the bilinear forms $\mathsf{A}_{\mathcal{M}}$ and $\mathsf{A}_{\mathcal{M}_{h}}$. Its proof is a straightforward adaptation of the proof of Dziuk and Elliott (2013, Lemma 4.7) to account for the diffusion matrix $\mathcal{D}$.
@@
Let $\Pi=I-\nu\nu^{T}$ (resp. $\Pi_{h}=I-\nu_{h}\nu_{h}^{T}$) be the orthogonal projection onto the tangent planes of $\mathcal{M}$ (resp. $\mathcal{M}_{h}$), and let $\mathcal{H}:\mathcal{M}\rightarrow\amsmathbb{R}^{(d+1)\times(d+1)}$ be the extended Weingarten map of $\mathcal{M}$ (cf. Dziuk and Elliott (2013, Definition 2.5)). Recall in particular that $\mathcal{H}(x)\nu(x)=0$ for any $x\in\mathcal{M}$, meaning in particular that $\mathcal{H}\Pi=\Pi\mathcal{H}=\mathcal{H}$. Finally, we introduce the map $\mathcal{Q}_{h}:\mathcal{M}\rightarrow\amsmathbb{R}^{(d+1)\times(d+1)}$ defined as
@@
$(\mathcal{D}^{-\ell}\nabla_{\mathcal{M}_{h}}u_{h})\cdot\nabla_{\mathcal{M}_{h}}\overline{v}_{h}$ $=\big{(}\mathcal{D}^{-\ell}\Pi_{h}(I-d_{s}\mathcal{H}^{-\ell})\Pi^{-\ell}(\nabla_{\mathcal{M}}u_{h}^{\ell})^{-\ell}\big{)}\cdot\big{(}\Pi_{h}(I-d_{s}\mathcal{H}^{-\ell})\Pi^{-\ell}(\nabla_{\mathcal{M}}\overline{v}_{h}^{\ell})^{-\ell}\big{)}$
$=\sigma\big{(}\mathcal{Q}_{h}^{-\ell}(\nabla_{\mathcal{M}}u_{h}^{\ell})^{-\ell}\big{)}\cdot(\nabla_{\mathcal{M}}\overline{v}_{h}^{\ell})^{-\ell},$
@@
Note that following Equations 44 and 9, $\mathsf{A}_{\mathcal{M}_{h}}^{\ell}$ satisfies for any $u_{h},v_{h}\in S_{h}$ the equality $\mathsf{A}_{\mathcal{M}_{h}}^{\ell}(u_{h}^{\ell},v_{h}^{\ell})=\mathsf{A}_{\mathcal{M}_{h}}(u_{h},v_{h})$. Therefore, for any $u_{h},v_{h}\in S_{h}$, we bound
@@
We now bound these two terms. Recall that Dziuk and Elliott (2013, Lemma 4.1) shows
@@
Since $\|d_{s}\|_{L^{\infty}(\mathcal{M}_{h})}\lesssim h^{2}$ by Dziuk and Elliott (2013, Lemma 4.1) and $\mathcal{H}$ is defined independently of $h$, we conclude that
@@
Using that $\big{\|}\Pi\Pi_{h}^{\ell}\Pi-\Pi\big{\|}\lesssim h^{2}$ by the proof of Dziuk and Elliott (2013, Lemma 4.1), we deduce that $\big{\|}\Pi\Pi_{h}^{\ell}\mathcal{D}\Pi_{h}^{\ell}\Pi-\mathcal{D}\big{\|}\lesssim h^{2}$. Injecting this inequality into Equation 50, and the resulting inequality into Equation 49, we conclude that
@@
$\big{|}\int_{\mathcal{M}}((\mathcal{Q}_{h}-\mathcal{D})$ $\nabla_{\mathcal{M}}u^{\ell}_{h})\cdot(\nabla_{\mathcal{M}}\overline{v}_{h}^{\ell})\,\mathrm{d}A\big{|}\leq\int_{\mathcal{M}}\|(\mathcal{Q}_{h}-\mathcal{D})\nabla_{\mathcal{M}}u^{\ell}_{h}\|\leavevmode\nobreak\ \|\nabla_{\mathcal{M}}{v_{h}}^{\ell}\|\,\mathrm{d}A$
$\lesssim\int_{\mathcal{M}}h^{2}\|\nabla_{\mathcal{M}}u^{\ell}_{h}\|\leavevmode\nobreak\ \|\nabla_{\mathcal{M}}{v_{h}}^{\ell}\|\,\mathrm{d}A\leq h^{2}\int_{\mathcal{M}}(\mu_{\min})^{-1}\|\nabla_{\mathcal{M}}u^{\ell}_{h}\|_{\mathcal{D}}\leavevmode\nobreak\ \|\nabla_{\mathcal{M}}{v_{h}}^{\ell}\|_{\mathcal{D}}\,\mathrm{d}A,$
@@
(51) $\big{|}\int_{\mathcal{M}}((\mathcal{Q}_{h}-\mathcal{D})\nabla_{\mathcal{M}}u^{\ell}_{h})\cdot(\nabla_{\mathcal{M}}\overline{v}_{h}^{\ell})\leavevmode\nobreak\ \,\mathrm{d}A\big{|}$ $\lesssim h^{2}\bigg{(}\int_{\mathcal{M}}\|\nabla_{\mathcal{M}}u^{\ell}_{h}\|_{\mathcal{D}}^{2}\leavevmode\nobreak\ \,\mathrm{d}A\bigg{)}^{1/2}\bigg{(}\int_{\mathcal{M}}\|\nabla_{\mathcal{M}}v^{\ell}_{h}\|_{\mathcal{D}}^{2}\leavevmode\nobreak\ \,\mathrm{d}A\bigg{)}^{1/2}.$
@@
B.2. Norm and error estimates of discrete operators
We now prove some estimates for the norm of shifted inverses of the operators $\mathcal{L}_{h}$ and $\mathsf{L}_{h}$, and for the error between inverses of these two operators. These results can be seen as extensions of the ones stated in Bonito et al. (2024, Lemma A.1).
@@ B.2. Norm and error estimates of discrete operators
Let $v_{h}\in S_{h}^{\ell}$ and $V_{h}\in S_{h}$ be arbitrary. Then, for all $z\in\Gamma$, for any $q\in[-1,1]$, and any $r\in(0,2)$,
(52) $\|\mathcal{L}_{h}^{1/2}(z-\mathcal{L}_{h})^{-1}v_{h}\|_{L^{2}(\mathcal{M})}\lesssim|z|^{-(1+q)/2}\|\mathcal{L}_{h}^{q/2}v_{h}\|_{L^{2}(\mathcal{M})},$
@@
We start with the proof of Equation 52. To this end, let $z\in\Gamma$ and $q\in[-1,1]$. For any $s\in[0,1]$ and $v_{h}\in S_{h}^{\ell}$, we expand
$\mathcal{L}_{h}^{s}v_{h}=\sum_{i=1}^{N_{h}}(\lambda_{i}^{h})^{s}(v_{h},e_{i}^{h})_{L^{2}(\mathcal{M})}e_{i}^{h},$
implying that
$\|\mathcal{L}_{h}^{s}v_{h}\|_{L^{2}(\mathcal{M})}^{2}=\sum_{i=1}^{N_{h}}|\lambda_{i}^{h}|^{2s}|(v_{h},e_{i}^{h})_{L^{2}(\mathcal{M})}|^{2}=\sum_{i=1}^{N_{h}}\left(|\lambda_{i}^{h}||(v_{h},e_{i}^{h})_{L^{2}(\mathcal{M})}|\right)^{2s}|(v_{h},e_{i}^{h})_{L^{2}(\mathcal{M})}|^{2(1-s)}.$
Hölder’s inequality yields
(56) $\begin{split}\|\mathcal{L}_{h}^{s}v_{h}\|_{L^{2}(\mathcal{M})}^{2}&\leq\left(\sum_{i=1}^{N_{h}}|\lambda_{i}^{h}|^{2}|(v_{h},e_{i}^{h})_{L^{2}(\mathcal{M})}|^{2}\right)^{s}\left(\sum_{i=1}^{N_{h}}|(v_{h},e_{i}^{h})_{L^{2}(\mathcal{M})}|^{2}\right)^{1-s}\\ &=\|\mathcal{L}_{h}v_{h}\|_{L^{2}(\mathcal{M})}^{2s}\|v_{h}\|_{L^{2}(\mathcal{M})}^{2(1-s)}.\end{split}$
Let $w_{h}=(z-\mathcal{L}_{h})^{-1}v_{h}$, then we obtain by Equation 20 that
$\|w_{h}\|_{L^{2}(\mathcal{M})}\lesssim|z|^{-1}\|v_{h}\|_{L^{2}(\mathcal{M})}.$
Hence, adding and subtracting $z(z-\mathcal{L}_{h})^{-1}$ yields
$\|\mathcal{L}_{h}w_{h}\|_{L^{2}(\mathcal{M})}$ $=\|z(z-\mathcal{L}_{h})^{-1}v_{h}-v_{h}\|_{L^{2}(\mathcal{M})}\leq|z|\|(z-\mathcal{L}_{h})^{-1}v_{h}\|_{L^{2}(\mathcal{M})}+\|v_{h}\|_{L^{2}(\mathcal{M})}$
$\lesssim 2\|v_{h}\|_{L^{2}(\mathcal{M})}.$
Combining this estimate with Equation 56, we get for any $s\in[0,1]$
(57) $\|\mathcal{L}_{h}^{s}w_{h}\|_{L^{2}(\mathcal{M})}\leq\|\mathcal{L}_{h}w_{h}\|_{L^{2}(\mathcal{M})}^{s}\|w_{h}\|_{L^{2}(\mathcal{M})}^{(1-s)}\leq|z|^{-(1-s)}\|v_{h}\|^{s}_{L^{2}(\mathcal{M})}\|v_{h}\|_{L^{2}(\mathcal{M})}^{1-s}=|z|^{s-1}\|v_{h}\|_{L^{2}(\mathcal{M})},$
yielding for any $q\in[-1,1]$ that
(58) $\begin{split}\|\mathcal{L}_{h}^{1/2}(z-\mathcal{L}_{h})^{-1}v_{h}\|_{L^{2}(\mathcal{M})}&=\|\mathcal{L}_{h}^{1-(1+q)/2}(z-\mathcal{L}_{h})^{-1}\mathcal{L}_{h}^{q/2}v_{h}\|_{L^{2}(\mathcal{M})}\\ &\lesssim|z|^{-(1+q)/2}\|\mathcal{L}_{h}^{q/2}v_{h}\|_{L^{2}(\mathcal{M})},\end{split}$
where we applied Equation 57 with $s=1-(1+q)/2$ in the last step. Hence, we retrieve Equation 52.
Next, to prove Equation 53, we first observe that the estimates in Equation 57 and Equation 58 carry over to the case when $\mathsf{L}_{h}$ is used instead of $\mathcal{L}_{h}$, as the proof in essence is a standard manipulation of a finite eigenexpansion. In particular, we obtain (by taking $q=1$ and $s=0$)
(59) $\|\mathsf{L}_{h}(z-\mathsf{L}_{h})^{-1}V_{h}\|_{L^{2}(\mathcal{M}_{h})}\lesssim|z|^{-1/2}\|\mathsf{L}_{h}^{1/2}V_{h}\|_{L^{2}(\mathcal{M}_{h})}.$
Now, Equation 59 combined with the equivalence of norms in Equation 19, yields that
$\|\mathsf{L}_{h}(z-\mathsf{L}_{h})^{-1}V_{h}\|_{L^{2}(\mathcal{M}_{h})}\lesssim|z|^{-1/2}\|V_{h}\|_{H^{1}(\mathcal{M}_{h})},$
which shows Equation 53.
@@
$\|(z-\mathcal{L}_{h})^{-1}v_{h}\|_{L^{2}(\mathcal{M})}=\|\mathcal{L}_{h}^{r/2}(z-\mathcal{L}_{h})^{-1}\mathcal{L}_{h}^{-r/2}v_{h}\|_{L^{2}(\mathcal{M})}\lesssim|z|^{r/2-1}\|\mathcal{L}_{h}^{-r/2}v_{h}\|_{L^{2}(\mathcal{M})}.$
@@
$(\mathsf{L}_{h}U_{h},W_{h})_{L^{2}(\mathcal{M}_{h})}=\mathsf{A}_{\mathcal{M}_{h}}(U_{h},W_{h})=(\mathsf{P}_{h}(\sigma v_{h}^{-\ell}),W_{h})_{L^{2}(\mathcal{M}_{h})}=(\sigma v_{h}^{-\ell},W_{h})_{L^{2}(\mathcal{M}_{h})}=(v_{h},W_{h}^{\ell})_{L^{2}(\mathcal{M})},$
@@
$|\mathsf{A}_{\mathcal{M}_{h}}(u_{h}^{-\ell}-U_{h},\Xi_{h})|$ $=|\mathsf{A}_{\mathcal{M}_{h}}(u_{h}^{-\ell},\Xi_{h})-\mathsf{A}_{\mathcal{M}_{h}}(U_{h},\Xi_{h})|=|\mathsf{A}_{\mathcal{M}_{h}}(u_{h}^{-\ell},\Xi_{h})-(v_{h},\Xi_{h}^{\ell})_{L^{2}(\mathcal{M})}|$
$=|\mathsf{A}_{\mathcal{M}_{h}}(u_{h}^{-\ell},\Xi_{h})-\mathsf{A}_{\mathcal{M}}(u_{h},\Xi_{h}^{\ell})|,$
@@
$\|\Xi_{h}^{\ell}\|_{H^{1}(\mathcal{M})}^{2}$ $\lesssim\mathsf{A}_{\mathcal{M}_{h}}(\Xi_{h},\Xi_{h})+\big{|}\mathsf{A}_{\mathcal{M}}(\Xi_{h}^{\ell},\Xi_{h}^{\ell})-\mathsf{A}_{\mathcal{M}_{h}}(\Xi_{h},\Xi_{h})\big{|}\lesssim\|\Xi_{h}\|_{H^{1}(\mathcal{M}_{h})}^{2}+h^{2}\|\Xi_{h}^{\ell}\|_{H^{1}(\mathcal{M})}^{2},$

### v2

@@ B.1. Geometric consistency estimate
The following geometric consistency estimate quantifies the error between the bilinear forms $\mathsf{A}_{\mathcal{M}}$ and $\mathsf{A}_{\mathcal{M}_{h}}$. Its proof is a straightforward adaptation of the proof of [11, Lemma 4.7] to account for the diffusion matrix $\mathcal{D}$.
@@
Let $\Pi=I-\nu\nu^{T}$ (resp. $\Pi_{h}=I-\nu_{h}\nu_{h}^{T}$) be the orthogonal projection onto the tangent planes of $\mathcal{M}$ (resp. $\mathcal{M}_{h}$), and let $\mathcal{H}:\mathcal{M}\rightarrow\amsmathbb{R}^{(d+1)\times(d+1)}$ be the extended Weingarten map of $\mathcal{M}$ (cf. [11, Definition 2.5]). Recall in particular that $\mathcal{H}(x)\nu(x)=0$ for any $x\in\mathcal{M}$, meaning in particular that $\mathcal{H}\Pi=\Pi\mathcal{H}=\mathcal{H}$. Finally, we introduce the map $\mathcal{Q}_{h}:\mathcal{M}\rightarrow\amsmathbb{R}^{(d+1)\times(d+1)}$ defined as
@@
$(\mathcal{D}^{-\ell}\nabla_{\mathcal{M}_{h}}{u_{h}}){\cdot}\nabla_{\mathcal{M}_{h}}{\overline{v}_{h}}$
$\qquad=\big{(}\mathcal{D}^{-\ell}\Pi_{h}(I-d_{s}\mathcal{H}^{-\ell})\Pi^{-\ell}(\nabla_{\mathcal{M}}u_{h}^{\ell})^{-\ell}\big{)}{\cdot}\big{(}\Pi_{h}(I-d_{s}\mathcal{H}^{-\ell})\Pi^{-\ell}(\nabla_{\mathcal{M}}\overline{v}_{h}^{\ell})^{-\ell}\big{)}$
$\qquad=\sigma\big{(}\mathcal{Q}_{h}^{-\ell}(\nabla_{\mathcal{M}}u_{h}^{\ell})^{-\ell}\big{)}\cdot(\nabla_{\mathcal{M}}\overline{v}_{h}^{\ell})^{-\ell},$
@@
Note that following Equations 49 and 9, $\mathsf{A}_{\mathcal{M}_{h}}^{\ell}$ satisfies for any $u_{h},v_{h}\in S_{h}$ the equality
$\mathsf{A}_{\mathcal{M}_{h}}^{\ell}(u_{h}^{\ell},v_{h}^{\ell})=\mathsf{A}_{\mathcal{M}_{h}}(u_{h},v_{h})$. Therefore, for any $u_{h},v_{h}\in S_{h}$, we bound
@@
We now bound these two terms. Recall that [11, Lemma 4.1] shows
@@
Since $\|d_{s}\|_{L^{\infty}(\mathcal{M}_{h})}\lesssim h^{2}$ by [11, Lemma 4.1] and $\mathcal{H}$ is defined independently of $h$, we conclude that
@@
Using that $\big{\|}\Pi\Pi_{h}^{\ell}\Pi-\Pi\big{\|}\lesssim h^{2}$ by the proof of [11, Lemma 4.1], we deduce that $\big{\|}\Pi\Pi_{h}^{\ell}\mathcal{D}\Pi_{h}^{\ell}\Pi-\mathcal{D}\big{\|}\lesssim h^{2}$. Injecting this inequality into Equation 55, and the resulting inequality into Equation 54, we conclude that
@@
$\left|\int_{\mathcal{M}}((\mathcal{Q}_{h}-\mathcal{D})\nabla_{\mathcal{M}}u^{\ell}_{h})\cdot(\nabla_{\mathcal{M}}\overline{v}_{h}^{\ell})\,\mathrm{d}A\right|\leq\int_{\mathcal{M}}\|(\mathcal{Q}_{h}-\mathcal{D})\nabla_{\mathcal{M}}u^{\ell}_{h}\|\leavevmode\nobreak\ \|\nabla_{\mathcal{M}}{v_{h}}^{\ell}\|\,\mathrm{d}A$
$\qquad\lesssim\int_{\mathcal{M}}h^{2}\|\nabla_{\mathcal{M}}u^{\ell}_{h}\|\leavevmode\nobreak\ \|\nabla_{\mathcal{M}}{v_{h}}^{\ell}\|\,\mathrm{d}A\leq h^{2}\int_{\mathcal{M}}(\mu_{\min})^{-1}\|\nabla_{\mathcal{M}}u^{\ell}_{h}\|_{\mathcal{D}}\leavevmode\nobreak\ \|\nabla_{\mathcal{M}}{v_{h}}^{\ell}\|_{\mathcal{D}}\,\mathrm{d}A,$
@@
(56) $\begin{split}&\left|\int_{\mathcal{M}}((\mathcal{Q}_{h}-\mathcal{D})\nabla_{\mathcal{M}}u^{\ell}_{h})\cdot(\nabla_{\mathcal{M}}\overline{v}_{h}^{\ell})\leavevmode\nobreak\ \,\mathrm{d}A\right|\\ &\qquad\lesssim h^{2}\bigg{(}\int_{\mathcal{M}}\|\nabla_{\mathcal{M}}u^{\ell}_{h}\|_{\mathcal{D}}^{2}\leavevmode\nobreak\ \,\mathrm{d}A\bigg{)}^{1/2}\bigg{(}\int_{\mathcal{M}}\|\nabla_{\mathcal{M}}v^{\ell}_{h}\|_{\mathcal{D}}^{2}\leavevmode\nobreak\ \,\mathrm{d}A\bigg{)}^{1/2}.\end{split}$
@@
B.2. Norm estimates
We start by introducing a few norm estimates involving the resolvents of $\mathcal{L}$ and $\mathcal{L}_{h}$. These results are straightforward extensions of the results introduced in [7, Lemmma 6.3], but adapted to the contour $\Gamma$ used to define functions of operators in this paper. We also recall a finite element error estimate which derives from a result in [7, Lemma 6.1].
@@ B.2. Norm estimates
For any $z\in\Gamma$, $s\in[0,1]$, $v\in L^{2}(\mathcal{M})$, and $v_{h}\in S_{h}^{\ell}$, it holds
(57) $\|\mathcal{L}^{s}(z-\mathcal{L})^{-1}v\|_{L^{2}(\mathcal{M})}$ $\lesssim|z|^{-(1-s)}\|v\|_{L^{2}(\mathcal{M})},$
(58) $\|\mathcal{L}_{h}^{s}(z-\mathcal{L}_{h})^{-1}v_{h}\|_{L^{2}(\mathcal{M})}$ $\lesssim|z|^{-(1-s)}\|v_{h}\|_{L^{2}(\mathcal{M})}.$
Besides, for any $\beta\in[0,1]$ and for any $\varphi\in L^{2}(\mathcal{M})$,
(59) $\|\mathcal{L}_{h}^{(1-\beta)/2}P_{h}(\mathcal{L}^{-1}-\mathcal{L}_{h}^{-1}P_{h})\varphi\|_{L^{2}(\mathcal{M})}\lesssim h^{2\beta}\|\mathcal{L}^{-(1-\beta)/2}\varphi\|_{L^{2}(\mathcal{M})}$
Proof.
We start with the proof of Equation 57. To this end, let $z\in\Gamma$ and $s\in[0,1]$. Let then $v\in L^{2}(\mathcal{M})$ and $w=(z-\mathcal{L})^{-1}v\in\mathcal{D}(\mathcal{L})$, where $\mathcal{D}(\cdot)$ denotes the domain of an operator. Since $\mathcal{L}$ is self-adjoint, we have $\mathcal{D}(\mathcal{L}^{s})=[L^{2}(\mathcal{M}),\mathcal{D}(\mathcal{L})]_{s}$ with isometry, where $[L^{2}(\mathcal{M}),\mathcal{D}(\mathcal{L})]_{s}$ the intermediate space between $L^{2}(\mathcal{M})$ and $\mathcal{D}(\mathcal{L})$ obtained by the complex interpolation method [28, Theorem 16.1]. In particular, we have
(60) $\|\mathcal{L}^{s}w\|_{L^{2}(\mathcal{M})}=\|w\|_{[L^{2}(\mathcal{M}),\mathcal{D}(\mathcal{L})]_{s}}\leq\|\mathcal{L}w\|_{L^{2}(\mathcal{M})}^{s}\|w\|_{L^{2}(\mathcal{M})}^{1-s}$
where the inequality derives from a classical result on interpolation spaces (see e.g., [28, Section 5.1]). On the one hand, we obtain by Equation 20 that
$\|w\|_{L^{2}(\mathcal{M})}=\|(z-\mathcal{L})^{-1}v\|_{L^{2}(\mathcal{M})}\lesssim|z|^{-1}\|v\|_{L^{2}(\mathcal{M})}.$
On the other hand, adding and subtracting $z(z-\mathcal{L})^{-1}v$ yields
$\|\mathcal{L}w\|_{L^{2}(\mathcal{M})}$ $=\|z(z{-}\mathcal{L})^{-1}v-v\|_{L^{2}(\mathcal{M})}\leq|z|\|(z{-}\mathcal{L})^{-1}v\|_{L^{2}(\mathcal{M})}+\|v\|_{L^{2}(\mathcal{M})}\lesssim 2\|v\|_{L^{2}(\mathcal{M})},$
where we once again use Equation 20 to derive the last inequality. Combining these two estimates with Equation 60, we get
$\|\mathcal{L}^{s}(z{-}\mathcal{L})^{-1}v\|_{L^{2}(\mathcal{M})}{=}\|\mathcal{L}^{s}w\|_{L^{2}(\mathcal{M})}\lesssim\|v\|^{s}_{L^{2}(\mathcal{M})}|z|^{-(1-s)}\|v\|_{L^{2}(\mathcal{M})}^{1-s}=|z|^{-(1-s)}\|v\|_{L^{2}(\mathcal{M})},$
hence proving Equation 57. As for Equation 58, it is proven using the same steps as Equation 57 but substituting $\mathcal{L}$ by $\mathcal{L}_{h}$, Equation 20 by Equation 21, and $\mathcal{D}(\mathcal{L})$ by $S_{h}^{\ell}$.
Finally, let us prove Equation 59. Let $\beta\in(0,1]$ and $\varphi\in L^{2}(\mathcal{M})$. Then, we have
$\|\mathcal{L}_{h}^{(1-\beta)/2}P_{h}(\mathcal{L}^{-1}-\mathcal{L}_{h}^{-1}P_{h})\varphi\|_{L^{2}(\mathcal{M})}$ $\lesssim\|\mathcal{L}^{(1-\beta)/2}P_{h}(\mathcal{L}^{-1}-\mathcal{L}_{h}^{-1}P_{h})\varphi\|_{L^{2}(\mathcal{M})}$
$\lesssim\|\mathcal{L}^{(1-\beta)/2}(\mathcal{L}^{-1}-\mathcal{L}_{h}^{-1}P_{h})\varphi\|_{L^{2}(\mathcal{M})}$
where we use [7, Lemma 5.2] to derive the first inequality, and [7, Lemma 5.1] to derive the second inequality. Finally, noting that $\mathcal{L}$ satisfies elliptic regularity for indices $\alpha\in(0,1]$ (cf. [7, Assumption 1]), we can apply [7, Lemma 6.1] with $\alpha=\beta$ and $s=(1-\beta)/2$ to conclude that $\|\mathcal{L}^{(1-\beta)/2}(\mathcal{L}^{-1}-\mathcal{L}_{h}^{-1}P_{h})\varphi\|_{L^{2}(\mathcal{M})}\lesssim h^{2\beta}\|\mathcal{L}^{-(1-\beta)/2}\varphi\|_{L^{2}(\mathcal{M})}$ and therefore
$\|\mathcal{L}_{h}^{(1-\beta)/2}P_{h}(\mathcal{L}^{-1}-\mathcal{L}_{h}^{-1}P_{h})\varphi\|_{L^{2}(\mathcal{M})}$ $\lesssim h^{2\beta}\|\mathcal{L}^{-(1-\beta)/2}\varphi\|_{L^{2}(\mathcal{M})}.$
This concludes the proof of Lemma B.2. ∎
We now prove some estimates for the norm of shifted inverses of the operators $\mathcal{L}_{h}$ and $\mathsf{L}_{h}$, and for the error between inverses of these two operators. These results can be seen as extensions of the ones stated in [6, Lemma A.1].
Lemma B.3.
Let $v\in L^{2}(\mathcal{M})$, $v_{h}\in S_{h}^{\ell}$ and $V_{h}\in S_{h}$ be arbitrary. Then, for all $z\in\Gamma$, for any $q\in[-1,1]$, and any $r\in(0,2)$, and any $s\in[0,1]$,
@@
To prove Equation 61, we first observe that the estimate in Equation 58 carry over to the case when $\mathsf{L}_{h}$ is used instead of $\mathcal{L}_{h}$, and $L^{2}(\mathcal{M})$ (resp. $S_{h}^{\ell}$) is replaced by its counterpart $L^{2}(\mathcal{M}_{h})$ (resp. $S_{h}$) on the polyhedral surface. Hence, for any $z\in\Gamma$, $V_{h}\in L^{2}(\mathcal{M}_{h})$, we have
$\|\mathsf{L}_{h}^{s}(z-\mathsf{L}_{h})^{-1}V_{h}\|_{L^{2}(\mathcal{M}_{h})}\lesssim|z|^{-(1-s)}\|V_{h}\|_{L^{2}(\mathcal{M}_{h})}$
In particular, we retrieve (by taking $s=1/2$)
(64) $\begin{split}\|\mathsf{L}_{h}(z-\mathsf{L}_{h})^{-1}V_{h}\|_{L^{2}(\mathcal{M}_{h})}&=\|\mathsf{L}_{h}^{1/2}(z-\mathsf{L}_{h})^{-1}\mathsf{L}_{h}^{1/2}V_{h}\|_{L^{2}(\mathcal{M}_{h})}\\ &\lesssim|z|^{-1/2}\|\mathsf{L}_{h}^{1/2}V_{h}\|_{L^{2}(\mathcal{M}_{h})}\lesssim|z|^{-1/2}\|V_{h}\|_{H^{1}(\mathcal{M}_{h})},\end{split}$
where we used the equivalence of norms (19) in the last inequality.
@@
$\|(z-\mathcal{L}_{h})^{-1}v_{h}\|_{L^{2}(\mathcal{M})}=\|\mathcal{L}_{h}^{r/2}(z-\mathcal{L}_{h})^{-1}\mathcal{L}_{h}^{-r/2}v_{h}\|_{L^{2}(\mathcal{M})}\lesssim|z|^{-(1-r/2)}\|\mathcal{L}_{h}^{-r/2}v_{h}\|_{L^{2}(\mathcal{M})}.$
@@
$(\mathsf{L}_{h}U_{h},W_{h})_{L^{2}(\mathcal{M}_{h})}$ $=\mathsf{A}_{\mathcal{M}_{h}}(U_{h},W_{h})=(\mathsf{P}_{h}(\sigma v_{h}^{-\ell}),W_{h})_{L^{2}(\mathcal{M}_{h})}$
$=(\sigma v_{h}^{-\ell},W_{h})_{L^{2}(\mathcal{M}_{h})}=(v_{h},W_{h}^{\ell})_{L^{2}(\mathcal{M})},$
@@
$|\mathsf{A}_{\mathcal{M}_{h}}(u_{h}^{-\ell}-U_{h},\Xi_{h})|$ $=|\mathsf{A}_{\mathcal{M}_{h}}(u_{h}^{-\ell},\Xi_{h})-\mathsf{A}_{\mathcal{M}_{h}}(U_{h},\Xi_{h})|$
$=|\mathsf{A}_{\mathcal{M}_{h}}(u_{h}^{-\ell},\Xi_{h})-(v_{h},\Xi_{h}^{\ell})_{L^{2}(\mathcal{M})}|=|\mathsf{A}_{\mathcal{M}_{h}}(u_{h}^{-\ell},\Xi_{h})-\mathsf{A}_{\mathcal{M}}(u_{h},\Xi_{h}^{\ell})|,$
@@
$\|\Xi_{h}^{\ell}\|_{H^{1}(\mathcal{M})}^{2}$ $\lesssim\mathsf{A}_{\mathcal{M}_{h}}(\Xi_{h},\Xi_{h})+\big{|}\mathsf{A}_{\mathcal{M}}(\Xi_{h}^{\ell},\Xi_{h}^{\ell})-\mathsf{A}_{\mathcal{M}_{h}}(\Xi_{h},\Xi_{h})\big{|}$
$\lesssim\|\Xi_{h}\|_{H^{1}(\mathcal{M}_{h})}^{2}+h^{2}\|\Xi_{h}^{\ell}\|_{H^{1}(\mathcal{M})}^{2},$